### v1

# Stacked conformal prediction

Abstract
We consider the conformalization of a stacked ensemble of predictive models, showing that the potentially simple form of the meta-learner at the top of the stack enables a procedure with manageable computational cost that achieves approximate marginal validity without requiring the use of a separate calibration sample. Empirical results indicate that the method compares favorably to a standard inductive alternative.
keywords:
Supervised learning, prediction sets, conformal prediction, stacking, exchangeability.
1 Introduction
In its most general transductive (full) form, conformal prediction (Vovk et al., 1999, 2005) creates a tension between desirable theoretical properties and computational complexity: while it always guarantees, for exchangeable data, the finite-sample marginal coverage of the nonparametric, model-agnostic prediction sets produced by the method, it may, in certain settings, require a highly intensive computational process, due to the necessity of repeatedly retraining the underlying predictive model. Inductive (split) conformal prediction (Papadopoulos et al., 2002; Lei et al., 2018) is an appealing and widely adopted variation that balances the computational cost through a trade-off: while preserving the statistical properties of the full conformal prediction procedure and requiring a single training process for the specified predictive model, it forces the available data to be split into training and calibration samples, which may result in neither building the most accurate predictive model nor obtaining the most efficient confidence assessment of the model’s generalization capacity. Several alternatives have been developed to keep the computational cost under control while making a more symmetric use of the available data, starting with cross-conformal prediction (Vovk, 2015), which inspired the development of the jackknife+ (Barber et al., 2021) and the nested conformal prediction framework (Gupta et al., 2022).
In this work, we develop a method for the conformalization of a stacked ensemble of predictive models (Wolpert, 1992; Breiman, 1996). This construction serves a dual purpose. First, it leverages the usual predictive gains of model stacking. Second, it takes advantage of the fact that the meta-learner at the top of the stack can be a simpler model, enabling a low-cost conformalization procedure that eliminates the need for a separate calibration set while producing prediction intervals with approximate marginal validity.
We proceed in the following manner. In Section 2, in a regression setting with exchangeable data, we construct a stack of predictive models that becomes totally symmetric by including the future observable pair in the training sample used to fit the base-learners at the first stack level. For this symmetric stack, the assumed exchangeability of the training sample is transferred to the second level of the stack. Consequently, for an arbitrary meta-learner placed at the top of the stack, we have the exact marginal validity of a random conformal prediction set associated with the future observable pair. Section 3 makes the stack construction feasible by moving the future observable pair out of the training sample. This breaks the distributional symmetry of the stack, but, under a stability assumption for the base stack models, the previously stated marginal validity property still holds approximately. In Section 4, we specialize the meta-learner to multiple linear regression to take advantage of the fact that for this simple model a full conformalization procedure can be carried out at low computational cost through standard rank-one matrix updates. We apply the resulting stacked conformal prediction algorithm to two datasets, comparing the stacked conformal prediction intervals to those produced by the inductive conformalized quantile regression method. We give pointers to open source software implementations of the developed methods in Section 5.
2 Symmetric stack
In a supervised learning setting, we have random pairs $\{(X_{i},Y_{i})\}_{i=1}^{n+1}$, for which the features vector $X_{i}\in\mathbb{R}^{d}$ and the response variable $Y_{i}\in\mathbb{R}$. We refer to $(X_{n+1},Y_{n+1})$ as the future observable pair. A permutation matrix is a square matrix in which each row and each column contains exactly one entry equal to $1$, and all other entries are equal to 00. Laying out our training sample — which for now includes the future observable pair — as the $(n+1)\times(d+1)$ random matrix
$T=\begin{pmatrix}X_{1,1}&X_{1,2}&\cdots&X_{1,d}&Y_{1}\\ X_{2,1}&X_{2,2}&\cdots&X_{2,d}&Y_{2}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ X_{n,1}&X_{n,2}&\cdots&X_{n,d}&Y_{n}\\ X_{n+1,1}&X_{n+1,2}&\cdots&X_{n+1,d}&Y_{n+1}\end{pmatrix},$
we make the assumption that the data are exchangeable, meaning that $T$ and $\Pi\,T$ have the same distribution for every permutation matrix $\Pi\in\mathbb{R}^{(n+1)\times(n+1)}$. We divide the training sample into $K\geq 2$ folds of size $t=(n+1)/K$, assuming that $n+1$ is divisible by $K$, by choosing a permutation matrix
$Q=\begin{pmatrix}Q_{1}\\ Q_{2}\\ \vdots\\ Q_{K}\end{pmatrix}\in\mathbb{R}^{(n+1)\times(n+1)},$
in which the matrix blocks $Q_{k}\in\mathbb{R}^{t\times(n+1)}$, for $k=1,\dots,K$. To identify the fold to which each training sample pair belongs, let $\phi:\{1,\dots,n+1\}\to\{1,\dots,K\}$ be a fold indicator function, defined by $\phi(i)=\sum_{k=1}^{K}k\cdot I(\sum_{\ell=1}^{t}(Q_{k})_{\ell,i}=1)$. Moreover, using the notation introduced for the $Q$ blocks, we define, for $k=1,\dots,K$, the $k$-th fold exclusion matrix
$Q_{\setminus k}=\begin{pmatrix}Q_{1}\\ \vdots\\ Q_{k-1}\\ Q_{k+1}\\ \vdots\\ Q_{K}\end{pmatrix}\in\mathbb{R}^{(n+1-t)\times(n+1)}.$
The stack is built from $M\geq 1$ base learning methods. For each training sample unit, the goal is to use its features to predict the response variable using models trained only on data folds that do not include the training sample unit under consideration. Formally, for $m=1,\dots,M$, we have prediction functions
$\hat{\mu}_{m}:\mathbb{R}^{(n+1-t)\times(d+1)}\times\mathbb{R}^{d}\to\mathbb{R},$
and we assume that each learning method treats its training data $S\in\mathbb{R}^{(n+1-t)\times(d+1)}$ symmetrically, so that
$\hat{\mu}_{m}(S,x)=\hat{\mu}_{m}(\Pi\,S,x),$
for every permutation matrix $\Pi\in\mathbb{R}^{(n+1-t)\times(n+1-t)}$ and $x\in\mathbb{R}^{d}$. When the learning method involves some form of randomization — such as the bootstrap process in Random Forests (Breiman, 2001), or the stochastic gradient descent optimization in Deep Neural Networks (Bishop and Bishop, 2023) — we assume that the seed of the underlying pseudo-random number generator is set using a symmetric hash function of the model’s training data $S$. The stack base-learners make predictions
$\hat{Z}_{i}=(\hat{Z}_{i,1},\dots,\hat{Z}_{i,M})\in\mathbb{R}^{M},$
in which $\hat{Z}_{i,m}=\hat{\mu}_{m}\!\left(Q_{\setminus\phi(i)}T,X_{i}\right)$, for $i=1,\dots,n+1$. This totally symmetric construction of the stack implies that the assumed exchangeability of the training sample $T$ is transferred to the next stack level.
Proposition 2.1.
The second-level random pairs $(\hat{Z}_{1},Y_{1}),\dots,(\hat{Z}_{n+1},Y_{n+1})$ are exchangeable for a symmetric stack.
The following result formalizes the idea that, for a generic meta-learner placed at the top of a symmetric stack, Proposition 2.1 and a standard conformal argument allow us to construct a random prediction set for the future observable pair with exact marginal validity. We denote the ceiling of $u\in\mathbb{R}$ by $\lceil u\rceil=\min\{z\in\mathbb{Z}:u\leq z\}$.
Proposition 2.2.
Let $\hat{\psi}_{n}^{(y)}$ be a meta-learner trained from $\{(\hat{Z}_{1},Y_{1}),\dots,(\hat{Z}_{n},Y_{n}),(\hat{Z}_{n+1},y)\}$, for $y\in\mathbb{R}$, and suppose the order of the pairs in this sample is irrelevant for the construction of $\hat{\psi}_{n}^{(y)}$. For a conformity function $\rho:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, define the conformity scores $R_{i}^{(y)}=\rho(Y_{i},\hat{\psi}^{(y)}_{n}(\hat{Z}_{i}))$, for $i=1,\dots,n+1$, and let $R^{(y)}_{(1)}\leq R^{(y)}_{(2)}\leq\dots\leq R^{(y)}_{(n)}$ denote the ordered conformity scores among $\{R^{(y)}_{1},R^{(y)}_{2}\dots,R^{(y)}_{n}\}$. Choosing a nominal miscoverage level $10<\alpha<$ 1, we have that
$P\!\left(Y_{n+1}\in C^{(\alpha)}_{n+1}\right)\geq 1-\alpha,$
in which we defined the random prediction set $C^{(\alpha)}_{n+1}=\left\{y\in\mathbb{R}:R_{n+1}^{(y)}\leq R_{(\lceil(1-\alpha)(n+1)\rceil)}^{(y)}\right\}$.
3 Feasible stack
The symmetric stack described in Section 2 is an oracle construct that cannot be implemented in practice, since at training time we do not know the observed value of the future response $Y_{n+1}$. A feasible stack is attained by removing the future observable pair $(X_{n+1},Y_{n+1})$ from the training sample. In doing so, we break the distributional symmetry of the stack, but if the predictions made by the stack base-learners stay stable after this single sample unit removal, we can argue that the marginal validity property stated in Proposition 2.2 still holds approximately, as will be corroborated by empirical results in Section 4.
Suppose that in this feasible stack the first level prediction $\hat{Z}_{n+1}$ is now made by base-learners trained on the whole training sample. Following the notations introduced in Proposition 2.2 for the symmetric stack, let $\tilde{R}_{n+1}^{(Y_{n+1})}$ and $\tilde{R}_{(\lceil(1-\alpha)(n+1)\rceil)}^{(Y_{n+1})}$ be the corresponding random conformity scores pertaining to the feasible stack, and let $\tilde{C}^{(\alpha)}_{n+1}$ be the associated random prediction set.
The following approximation result shows how the marginal validity is impacted by the broken distributional symmetry of the feasible stack: if the base-learners are stable in the sense that we can control the probability of the corresponding conformity scores on both the symmetric and the feasible stack differing too much, the marginal validity property still approximately for the feasible stack.
Proposition 3.1.
Given $\epsilon>0$, if there is a $\delta=\delta(\epsilon)>0$, such that
$P\!\left(\max\left\{\left|R_{n+1}^{(Y_{n+1})}-\tilde{R}_{n+1}^{(Y_{n+1})}\right|,\left|R_{(\lceil(1-\alpha)(n+1)\rceil)}^{(Y_{n+1})}-\tilde{R}_{(\lceil(1-\alpha)(n+1)\rceil)}^{(Y_{n+1})}\right|\right\}<\epsilon/2\right)\geq 1-\delta,$
then
$P\!\left(Y_{n+1}\in\tilde{C}^{(\alpha)}_{n+1}\right)\geq 1-\alpha-\delta-h(\epsilon),$
in which $h(\epsilon)=P\!\left(\tilde{R}_{(\lceil(1-\alpha)(n+1)\rceil)}^{(Y_{n+1})}-\epsilon<\tilde{R}_{n+1}^{(Y_{n+1})}\leq\tilde{R}_{(\lceil(1-\alpha)(n+1)\rceil)}^{(Y_{n+1})}\right)$.
4 Specializing the meta-learner
The feasible stack described in Section 3 is useful only if we can make the computational cost involved in the conformalization of the meta-learner manageable. It has long been known (Vovk et al., 1999) that full conformalization of a multiple linear regression model can be carried out efficiently through classical rank-one matrix updates made with the help of the Sherman-Morrison formula (Golub and Van Loan, 2013).
Using multiple linear regression as the meta-learner of the feasible stack, we implemented the stacked conformal prediction procedure for two examples. The California housing dataset (Pace and Barry, 1997) comprises information on 20,640 census tracts in California, drawn from the 1990 U.S. Census. It includes eight predictors and the response variable is the median house value in US dollars for each tract. The Ames housing dataset (De Cock, 2011) contains information on 2,930 houses sold in Ames, Iowa from 2006 to 2010. For the Ames dataset, we have eighty predictors and the response variable is the sale price of the house in US dollars.
We used Random Forests (Breiman, 2001) and CatBoost (Prokhorenkova et al., 2018) as the base-learners of the stack. For both datasets, we used 70% of the available data as our training sample, and generated prediction sets for the remaining 30% test sample units. For comparison, we used conformalized quantile regression (CQR) (Romano et al., 2019) as an inductive alternative. The Quantile Random Forest algorithm (Meinshausen, 2006) was used in the implementation of CQR. For CQR, the calibration samples sizes were chosen to be 2,000 and 500, for the California and Ames datasets, respectively. In both examples, we used a 90% nominal miscoverage level. Figures 1 and 2 show the prediction intervals for fifty randomly chosen test sample units. Table 1 summarizes the results, showing that the stacked conformal prediction intervals exhibited appropriate empirical coverage, with shorter median length, when compared to the CQR results.
Figure 1: Prediction intervals for fifty test sample units in the California Housing dataset.
Figure 2: Prediction intervals for fifty test sample units in the Ames Housing dataset.
Table 1: Prediction interval length summaries (US dollars) and empirical coverage.
Dataset Method 1st quartile Median 2nd quartile Coverage
California Stacked CP 96,927 119,003 147,988 89.9%
CQR 104,433 156,600 231,500 90.8%
Ames Stacked CP 41,652 59,478 88,088 91.1%
CQR 44,530 68,960 108,839 90.2%
5 Open source implementations
R (R Core Team, 2024) and Python (Python Core Team, 2023) codes for the full reproduction of the examples in the paper are available at:
https://github.com/paulocmarquesf/stacked_cp
[…]
Paulo C. Marques F. receives support from FAPESP (Fundação de Amparo à Pesquisa do Estado de São Paulo) through project 2023/02538-0.